OpenStudy (anonymous):

you have 3/6 cup of cornmeal in your cupboard . A recipe for muffins calls for 1/4 cup of cornmeal .How much cornmeal would you have left if you made the muffins ?

OpenStudy (anonymous):

3/6-1/4=1/2-1/4=1/4 just subtract what u use from what you have

OpenStudy (anonymous):

this goes for every problem like this one ?

OpenStudy (anonymous):

yes